A Gronwall lemma for the weak solution of the system of Navier and Stokes

(pointwise selecting the solutions with minimal dissipation of energy)

Reinhold Böhme Ruhr-Universität Bochum Fakultät für Mathematik Universitätsstraße 150 Gebäude NA 5/72 D44780 Bochum Germany

            Reinhold.Boehme@ruhr-uni-bochum.de